OpenStudy (anonymous):

Using laplace trasform solve : dx/dt+3x=0

hartnn (hartnn):

have u tried it ? U know L[dx/dt] = ?

OpenStudy (anonymous):

i know the ans is 2e^-3t by solving it normally ... but i don/t know the lplace of dx/dt?

OpenStudy (anonymous):

i meant laplace lol

hartnn (hartnn):

laplace of derivative is L[dx/dt] = sF(s)-F(0)

hartnn (hartnn):

where F(s) is laplace of f(x)

hartnn (hartnn):

L[df(x)/dt] = sF(s)-F(0)

OpenStudy (anonymous):

i missed the intial value x(0)=2.. still confused?

hartnn (hartnn):

i am doing some typos here, sorry, L[df(x)/dt] = sF(s)-f(0) here your f(x) is x .

hartnn (hartnn):

f(0)=x(0)

OpenStudy (anonymous):

right so is it sF(s)-f(0) +3/s^2

hartnn (hartnn):

yes, put f(0) as 2

OpenStudy (anonymous):

if f(0)=2 then what about sF(s)?

hartnn (hartnn):

isolate F(s) i think u know that to solve it, we need to take inverse laplace transform also, by first finding F(s) in terms of s

OpenStudy (anonymous):

2/s+2) = 2e^-2t correct or not?

hartnn (hartnn):

3/(s+2) 3e^{-2t}

OpenStudy (anonymous):

if you solve the normal way you get 2e^-3t ? how come it is a diffrent ans?

OpenStudy (anonymous):

dx/dt = -3x (-1/3)(dx/x) = dt dx/x = -3dt ln x = -3t + C x = e^(-3t+C) x = Ce^(-3t) [The sloppiness I mentioned] 2 = Ce^0 C = 2 x(t) = 2e^(-3t)

hartnn (hartnn):

dx/dt+3x=0 sF(s)- f(0)+3F(s)=0 sF(s)- 2+3F(s)=0 (s+3)F(s) = 2 F(s) = 2/(s+3) f(x) = 2e^{-3t} i think i should sleep.....

hartnn (hartnn):

ask if u didn't get any of the step above.....

OpenStudy (anonymous):

i under now ... lol im new to the topic thats why im so confused !

hartnn (hartnn):

with practice, this topic becomes simpler ....

OpenStudy (anonymous):

understand i mean rofl

OpenStudy (anonymous):

how would i get about solving the laplace of d^2x/dt^2

OpenStudy (anonymous):

Like is their a table for this ?

hartnn (hartnn):

yes, there is table ... without initial conditions, L(d^2 f(x)/dt^2) = s^2 F(s)

OpenStudy (anonymous):

oh i see... now!

hartnn (hartnn):

any more doubts ?

OpenStudy (anonymous):

how did you get 3F(s) for the last Question?

OpenStudy (anonymous):

im just looking back now .. and noticed...

hartnn (hartnn):

laplace of x(t) = F(s)

OpenStudy (anonymous):

i thought 3x= 3/s^2 according to tables?

hartnn (hartnn):

its not x, its x(t), right ?

OpenStudy (anonymous):

ohh damn your right!..

OpenStudy (anonymous):

any chance of working through another question?

hartnn (hartnn):

any more questions ?

OpenStudy (anonymous):

yes , im working on dx/dt+2x=2te^-2t ; x(0)=3

hartnn (hartnn):

so what is laplace of dx/dt+2x ?

OpenStudy (anonymous):

sF(s)- f(0)+2F(s)=> 2te^-2t how do i use splace for this?

hartnn (hartnn):

asking laplace for this 2te^-2t, right ?

OpenStudy (anonymous):

yes .. do i use tables

hartnn (hartnn):

yes, from that what is L{t^n f(t)} = ?

OpenStudy (anonymous):

2/s^3 and 1/s+2 multiplied together i guess?

hartnn (hartnn):

no, laplace of product doesn't equal product of laplace \(\large L(tf(x))=-d/ds(F(s))\)

hartnn (hartnn):

so first find L(e^{-2t})

OpenStudy (anonymous):

so how im i supposed to solve 2te^-2t , if tables can't be used?

hartnn (hartnn):

use table to find L(e^{-2t}) first

OpenStudy (anonymous):

L(e^{-2t})= 1/s+2

hartnn (hartnn):

yes, so now L{t e^{-2t}} will be -d/ds(1/s+2) got this ?

OpenStudy (anonymous):

oh so you find the laplace of 2t now ?

OpenStudy (anonymous):

where is -d/ds coming from is that a standard result?

hartnn (hartnn):

no! according to rule \(\large L(tf(x))=-d/ds(F(s))\) you first find the laplace of the function that is mul.tiplied by t then take its derivative(of laplace) w.r.t s

hartnn (hartnn):

yes, its standard result

OpenStudy (anonymous):

right, i have never been taught that.... first time seeing this! WOW! lol

OpenStudy (anonymous):

so does that mean the laplace of 2 is solved separately?

hartnn (hartnn):

2 is constant, it can be taken out of laplace, just like in derivatives or integration

OpenStudy (anonymous):

can't you work it out like this L(2t)*L(e^-2t)

hartnn (hartnn):

no, as i said , laplace of product is not product of laplace \(L(f(x)g(x))\ne L(f(x))L(g(x))\)

OpenStudy (anonymous):

right thats wierd beacuse i was taught that it is linearly propotional and can be solved this way.. my lecturer is quiet confusing i guess!

hartnn (hartnn):

so now can u find L{t e^{-2t}}

OpenStudy (anonymous):

sF(s)- 3+2F(s)=-d/ds(1/s+2) ... how can i proceed forward when i have -d/ds?

hartnn (hartnn):

d/ds is just the first derivative w.r.t x so what is d/ds(1/(s+2)) = ?

OpenStudy (anonymous):

i don't get why you suppose to diffrentiate?

hartnn (hartnn):

because the formula says so :P

OpenStudy (anonymous):

-1/s^2 i think...

hartnn (hartnn):

d/ds(1/(s+2)) = -1/(s+2)^2

OpenStudy (anonymous):

yep i made mistake.. so i got (s+2) f(s)= 1/(s+2)^2+3 is that correct...

hartnn (hartnn):

yes, isolate F(s)

OpenStudy (anonymous):

f(s)= 1/(s+2)^2 +3/(s+3) F(x)= t^2e^-2t +3e^-2t

hartnn (hartnn):

(s+2) F(s)= 1/(s+2)^2+3 F(s) = 1/(s+2)^3 +3/(s+3) isn't it ?

OpenStudy (anonymous):

1/(s+2)^3 i think its 1/(s+2)^2

hartnn (hartnn):

(s+2) f(s)= 1/(s+2)^2+3 dividing s+2 throughout the equation, what u get ?

OpenStudy (anonymous):

yh your right! , but looking at tables their is no solution to 1/(s+2)^3

OpenStudy (anonymous):

finally figured it out, it is 1/(s+2)^3....

hartnn (hartnn):

1/s^3 ?

hartnn (hartnn):

for 1/(s+2), it is ?

OpenStudy (anonymous):

yes the power of t is added in to order to match 1/(s+a)^2

hartnn (hartnn):

thats correct, so what u get for 1/(s+2)^3 ?

OpenStudy (anonymous):

F(x)= t^2e^-2t +3e^-2t

OpenStudy (anonymous):

final ans

hartnn (hartnn):

first term will be divided by 2, isn't it ?

OpenStudy (anonymous):

f(s)= 1/(s+2)^3 +3/(s+3) F(x)= t^2e^-2t +3e^-2t

hartnn (hartnn):

all that is correct, i was saying that it should be F(x)=(1/2) t^2e^-2t +3e^-2t

OpenStudy (anonymous):

where is the 1/2 coming from?

hartnn (hartnn):

what is laplace inverse of 1/s^3 ?

OpenStudy (anonymous):

according to tables it seems right?

hartnn (hartnn):

according to table laplace inverse of 1/s^3 = t^2/2 isn't it ?

OpenStudy (anonymous):

t^2

hartnn (hartnn):

L(t^n)=n!/(s^(n+1))

OpenStudy (anonymous):

Im not sure ...

hartnn (hartnn):

isn't that in the list ?

OpenStudy (anonymous):

i did this f(t)=t^m , and according to f(s)=m!/s^m+1 M=2 ??? or m=1

hartnn (hartnn):

m=2, because power is 2

OpenStudy (anonymous):

well thats like saying 2x1/s^3 which is incorrect ... when it is 1/(s+2)^3

hartnn (hartnn):

i meant laplace inverse of 1/x^3 = t^2/ 2

OpenStudy (anonymous):

i think it is that... not sure 100%

hartnn (hartnn):

i am pretty sure about that . laplace inverse of 1/x^3 = t^2/ 2

OpenStudy (anonymous):

if your sure then i guess it must be correct !

hartnn (hartnn):

so final answer is F(x)=(1/2) t^2e^-2t +3e^-2t